OpenStudy (roxy.girl):

Help me.. -4.7 = 3y + 1.3

rishavraj (rishavraj):

@Roxy.girl wht would be the first step ??

OpenStudy (roxy.girl):

So you add negative 3y right.. @Qwertty123

OpenStudy (qwertty123):

Yes XD

rishavraj (rishavraj):

no need to u see u can ecen start by adding -1.3 on both sides @Roxy.girl

OpenStudy (qwertty123):

Reorder the terms: -4.7 = 1.3 + 3y Solving -4.7 = 1.3 + 3y Move all terms containing y to the left, all other terms to the right. Add '-3y' to each side of the equation. -4.7 + -3y = 1.3 + 3y + -3y

rishavraj (rishavraj):

it would be like - 6 = 3y and then u know wht to do

OpenStudy (roxy.girl):

huh?

rishavraj (rishavraj):

@Roxy.girl see it wont really make any difference tht u move the variable on right side....

OpenStudy (qwertty123):

Combine like terms: 3y + -3y = 0 -4.7 + -3y = 1.3 + 0 -4.7 + -3y = 1.3 Then: Add '4.7' to each side of the equation. -4.7 + 4.7 + -3y = 1.3 + 4.7

rishavraj (rishavraj):

like -4.7 = 3y + 3 -6 = 3y -2 = y see no difference :)) @Roxy.girl

OpenStudy (roxy.girl):

I am getting confused lol

OpenStudy (roxy.girl):

Now you would divide by negative 3? Right? @Qwertty123

OpenStudy (qwertty123):

Then Combine like terms: -4.7 + 4.7 = 0.0 0.0 + -3y = 1.3 + 4.7 -3y = 1.3 + 4.7 Combine like terms: 1.3 + 4.7 = 6 -3y = 6 Divide each side by '-3'. so yes your correct! :3

rishavraj (rishavraj):

@Roxy.girl the only difference is tht ur first step was subtracting -3y on both sides so I told tht its not a necessary step ...u can proceed without tht step

OpenStudy (roxy.girl):

oh lol okay

OpenStudy (roxy.girl):

So it equals -2. @Qwertty123

OpenStudy (qwertty123):

y = -2 yes :)

rishavraj (rishavraj):

yup I did mention tht in one of my replies -_-

OpenStudy (roxy.girl):

oh lol whoops